# PaddleOCR-VL: Boosting Multilingual Document Parsing via a 0.9B Ultra-Compact Vision-Language Model

### v3

@@ 4.1 Page-level Evaluation
Table 3: Comprehensive evaluation of document parsing on OmniDocBench v1.0. Results are reported by OmniDocBench [16] unless MinerU2.5 and Ours.
@@
olmOCR-Bench [12] includes 1,402 PDF documents and 7,010 test cases, addressing diverse document types and extraction challenges. It offers a detailed evaluation framework for PDF content extraction by assessing tools and models through simple, clear, and machine-verifiable unit tests. This approach avoids biased evaluations and soft metric comparisons, allowing for the detection of subtle but significant extraction errors.
@@
GOT [21] 48.3 ± 1.1 52.7 52.0 0.2 22.1 93.6 42.0 29.9 94.0
Gemini Flash 2 (No Anchor) [48] 57.8 ± 1.1 32.1 56.3 61.4 27.8 48.0 58.7 84.4 94.0
MinerU-pipeline [9] 61.5 ± 1.1 75.4 47.4 60.9 17.3 96.6 59.0 39.1 96.6
Gemini Flash 2 (Anchored) [48] 63.8 ± 1.2 54.5 56.1 72.1 34.2 64.7 61.5 71.5 95.6
Nanonets-OCR-s [51] 64.5 ± 1.1 67.0 68.6 77.7 39.5 40.7 69.9 53.4 99.3
Qwen2.5-VL-7B (No Anchor) [24] 65.5 ± 1.2 63.1 65.7 67.3 38.6 73.6 68.3 49.1 98.3
GPT-4o (No Anchor) [7] 68.9 ± 1.1 51.5 75.5 69.1 40.9 94.2 68.9 54.1 96.7
GPT-4o (Anchored) [7] 69.9 ± 1.1 53.5 74.5 70.0 40.7 93.8 69.3 60.6 96.8
Marker-1.8.2 [45] 70.1 ± 1.1 76.0 57.9 57.6 27.8 84.9 72.9 84.6 99.1
olmOCR v0.1.75 (No Anchor) [12] 74.7 ± 1.1 71.5 71.4 71.4 42.8 94.1 77.7 71.0 97.8
olmOCR v0.1.75 (Anchored) [12] 75.5 ± 1.0 74.9 71.2 71.0 42.2 94.5 78.3 73.3 98.3
MonkeyOCR-pro-3B [1] 75.8 ± 1.0 83.8 68.8 74.6 36.1 91.2 76.6 80.1 95.3
MinerU2.5 [2] 77.5 ± 1.0 81.1 74.0 85.1 33.8 96.3 65.5 89.8 94.4
dots.ocr [52] 79.1 ± 1.0 82.1 64.2 88.3 40.9 94.1 82.4 81.2 99.5
@@
Table 4: Comprehensive evaluation of document parsing on olmOCR-Bench. Results are reported by olmOCR-Bench [12] unless MinerU2.5 and Ours.
@@ 4.2.1 Text Recognition
Qwen2.5-VL-72B [24] 0.054 0.023 0.061 0.084 0.195 0.032 0.056 0.118 0.040
MonkeyOCR-pro-3B [1] 0.058 0.021 0.064 0.096 0.116 0.023 0.058 0.124 0.052
MinerU2.5 [2] 0.195 0.089 0.111 0.234 0.194 0.147 0.056 0.142 0.094
Dolphin [3] 0.237 0.095 0.135 0.347 0.248 0.233 0.121 0.309 0.213
@@
Qwen2.5-VL-72B [24] 0.405 0.056 0.389 0.165 0.194 0.758 0.164 0.220 0.021 0.181
Dolphin [3] 0.682 0.699 0.912 0.691 0.709 0.832 0.818 0.549 0.037 0.309
MonkeyOCR-pro-3B [1] 0.601 0.182 0.921 0.449 0.876 0.909 0.896 0.387 0.036 0.262
MinerU2.5 [2] 0.978 0.917 0.957 0.661 0.880 0.937 0.915 0.832 0.063 0.588
@@
Dolphin [3] 0.236 0.145 0.074 0.025 0.095 0.218 0.113 0.183 0.092 0.190 0.202 0.225 0.230
MonkeyOCR-pro-3B [1] 0.253 0.071 0.048 0.023 0.295 0.529 0.144 0.165 0.063 0.086 0.110 0.184 0.263
Qwen2.5-VL-72B [24] 0.188 0.047 0.037 0.018 0.100 0.387 0.122 0.186 0.034 0.090 0.041 0.134 0.220
MinerU2.5 [2] 0.370 0.088 0.041 0.023 0.232 0.950 0.179 0.256 0.048 0.962 0.097 0.174 0.337
@@
This is a line and paragraph levels handwritten evaluation dataset designed for comprehensive handwriting recognition assessment. It contains 400 samples, evenly divided into four subsets of 100 images each. The dataset covers both real and synthetic handwriting in Chinese and English. Real samples are collected from established handwriting datasets such as CASIA-HWDB [29], GNHK [59], and BRUSH [60], while synthetic samples are generated to simulate diverse writing styles, character densities, and layouts. The benchmark aims to provide balanced and fine-grained evaluation for handwritten text recognition across different scripts and writing conditions.
@@
InternVL2.5-4B [57] 0.197 0.240 0.661 0.741 0.674 0.754 0.655 0.734 0.406 0.473 0.652 0.687
MiniCPM-V2.6-8B [61] 0.147 0.175 0.727 0.810 0.747 0.811 0.714 0.812 0.443 0.583 0.727 0.774
Qwen2-VL-7B [62] 0.127 0.113 0.760 0.881 0.773 0.884 0.754 0.884 0.490 0.666 0.756 0.859
GOT [21] 0.616 0.402 0.283 0.568 0.309 0.618 0.273 0.544 0.151 0.295 0.255 0.492
PaddleOCR [10] 0.418 0.325 0.237 0.664 0.232 0.646 0.263 0.700 0.069 0.431 0.236 0.648
@@
Ocean-OCR [63] 0.145 0.106 0.774 0.885 0.780 0.912 0.782 0.862 0.532 0.736 0.772 0.885
MinerU2.5 [2] 0.238 0.356 0.558 0.619 0.547 0.623 0.574 0.622 0.344 0.489 0.553 0.601
@@
Table 7: Comparison of performance on English(EN) and Chinese(ZH) OCR for handwritten recognition on Ocean-OCR-Bench. Results are reported by Ocean-OCR [63] unless MinerU2.5 and Ours.
@@
For table recognition, we utilize two benchmarks to validate the effectiveness of PaddleOCR-VL-0.9B based on TEDS [41] and Edit Distance.
@@
dots.ocr [52] 0.8194 0.8442 0.1508
MinerU2.5 [2] 0.9005 0.9539 0.0693
@@
MinerU2-VLM [14] 0.8286 0.8730 0.8757 0.9088
MonkeyOCR [1] 0.7396 0.7824 0.8174 0.8537
Nanonets-OCR-s [51] 0.7824 0.8190 0.8377 0.8692
OCRFlux-3B [49] 0.7741 0.8071 0.8238 0.8617
Qwen2.5-VL-3B [24] 0.7398 0.7765 0.8132 0.8701
Qwen2.5-VL-7B [24] 0.7549 0.7926 0.8251 0.8819
Qwen2.5-VL-72B [24] 0.7762 0.8361 0.843 0.8987
dots.ocr [52] 0.7547 0.7914 0.8047 0.8361
MinerU2.5 [2] 0.8469 0.8955 0.8896 0.9239
@@
For formula recognition, we validate the effectiveness our model based on the Character Detection Matching (CDM) [64] metric on OmniDocBench-Formula-block and In-house-Formula datasets.
@@
dots.ocr [52] 0.4641 0.4868 0.4414
MinerU2-VLM [14] 0.8286 0.9616 0.6956
MonkeyOCR-pro-1.2B [1] 0.8531 0.9642 0.7419
MonkeyOCR-3B [1] 0.8621 0.9718 0.7524
Qwen2.5-VL-72B [24] 0.8747 0.9574 0.7920
MinerU2.5 [2] 0.9187 0.9751 0.8623
@@
Table 10: Comparison of OmniDocBench v1.5 Formula-block Performance. Due to dots.ocr [52] easily recognizing cropped formulas as images, the score is relatively low.
@@
dots.ocr [52] 0.6737 0.8066 0.5408
MinerU2-VLM [14] 0.9237 0.9764 0.8709
MonkeyOCR-pro-1.2B [1] 0.9537 0.9656 0.9417
MonkeyOCR-3B [1] 0.9566 0.9761 0.9371
Qwen2.5-VL-72B [24] 0.9412 0.9519 0.9304
MinerU2.5 [2] 0.9770 0.9832 0.9708
@@
Table 11: Comparison of In-house-Formula Performance. Due to dots.ocr [52] easily recognizing cropped formulas as images, the score is relatively low.
@@
For chart recognition, considering the limitations in dataset size, the imbalanced distribution of chart categories, and the poor annotation quality of publicly available test sets, we only utilize a in-house benchmark to validate the effectiveness of PaddleOCR-VL-0.9B based on the RMS-F1 [42] score metric. As shown in Table 12, the proposed PaddleOCR-VL not only outperforms expert OCR VLMs but also surpasses some 72B-level multimodal language models.
@@
TinyChart [65] 0.2159 0.4726 0.0876
GOT [21] 0.3160 0.1100 0.4190
OneChart [66] 0.3716 0.1384 0.4882
Qwen2.5-VL-3B [24] 0.5942 0.5619 0.6103
Qwen2.5-VL-7B [24] 0.6821 0.5876 0.7293
Qwen2.5-VL-72B [24] 0.7300 0.6972 0.7464
PP-StructureV3 [10] 0.8060 0.7963 0.8109
@@ 4.3 Inference Performance
To improve the inference performance of PaddleOCR-VL, we introduce multi-threading asynchronous execution into the inference workflow. The process is divided into three main stages—data loading (e.g., rendering PDF pages as images), layout model processing, and VLM inference—each running in a separate thread. Data is transferred between adjacent stages via queues, enabling concurrent execution for higher efficiency. In particular, for VLM inference, batch processing is only triggered when either the number of items in the queue reaches a predefined threshold or the waiting time for queued data exceeds a specified limit. This design allows blocks across different pages to be aggregated and processed together, thereby maximizing parallelism, especially when handling large volumes of files. We further deploy PaddleOCR-VL-0.9B on high-throughput inference and serving engines [67, 68, 69], tuning parameters like max-num-batched-tokens and gpu-memory-utilization to balance inference throughput with GPU memory consumption.
@@
MonkeyOCR-pro-1.2B [1] vLLM (v0.10.2) 1456.4 0.6730 1120.3
dots.ocr [52] vLLM (v0.10.2) 2784.6 0.3522 532.9
MinerU2.5 [2] vLLM (v0.10.2) 927.3 1.0574 1647.9
@@ A.2 Table
As shown in Figure A2, we constructed a large-scale dataset of over $5$ million high-quality image-table pairs. Our dataset construction employs three key strategies: automatic data annotation, potential annotation mining, and high-quality data synthesis. For coding efficiency, we adopt OTSL [28] as the model’s target format instead of conventional HTML. The main dataset construction process is as follows:
@@
    Automatic Data Annotation: To enhance the performance of PaddleOCR-VL in table recognition, we built a large-scale, diverse dataset covering various languages, border styles, and table types. Tables are first located using PP-StructureV3 [10]. For unlabeled images, we employed a multi-stage annotation pipeline: ERNIE-4.5-VL [5] first generates pseudo-labels, which are then validated by a ERNIE-4.5-VL-28B-A3B [5] as discriminative model. Rejected annotations are refined using DianJin-OCR-R1 [70] (for tools, we use ERNIE-4.5-VL and PP-StructureV3 [10]). Finally, all annotations undergo rigorous rule-based verification, including n-gram analysis and HTML validation, to ensure only high-quality samples are used for training.

### v4

@@ 4.1 Page-level Evaluation
Table 3: Comprehensive evaluation of document parsing on OmniDocBench v1.0. Results are reported by OmniDocBench [ouyang2025omnidocbench] unless MinerU2.5 and Ours.
@@
olmOCR-Bench [poznanski2025olmocr] includes 1,402 PDF documents and 7,010 test cases, addressing diverse document types and extraction challenges. It offers a detailed evaluation framework for PDF content extraction by assessing tools and models through simple, clear, and machine-verifiable unit tests. This approach avoids biased evaluations and soft metric comparisons, allowing for the detection of subtle but significant extraction errors.
@@
GOT [wei2024general] 48.3 ± 1.1 52.7 52.0 0.2 22.1 93.6 42.0 29.9 94.0
Gemini Flash 2 (No Anchor) [gemini25] 57.8 ± 1.1 32.1 56.3 61.4 27.8 48.0 58.7 84.4 94.0
MinerU-pipeline [wang2024mineru] 61.5 ± 1.1 75.4 47.4 60.9 17.3 96.6 59.0 39.1 96.6
Gemini Flash 2 (Anchored) [gemini25] 63.8 ± 1.2 54.5 56.1 72.1 34.2 64.7 61.5 71.5 95.6
Nanonets-OCR-s [Nanonets-OCR-S] 64.5 ± 1.1 67.0 68.6 77.7 39.5 40.7 69.9 53.4 99.3
Qwen2.5-VL-7B (No Anchor) [bai2025qwen2] 65.5 ± 1.2 63.1 65.7 67.3 38.6 73.6 68.3 49.1 98.3
GPT-4o (No Anchor) [achiam2023gpt] 68.9 ± 1.1 51.5 75.5 69.1 40.9 94.2 68.9 54.1 96.7
GPT-4o (Anchored) [achiam2023gpt] 69.9 ± 1.1 53.5 74.5 70.0 40.7 93.8 69.3 60.6 96.8
Marker-1.8.2 [vik2024marker] 70.1 ± 1.1 76.0 57.9 57.6 27.8 84.9 72.9 84.6 99.1
olmOCR v0.1.75 (No Anchor) [poznanski2025olmocr] 74.7 ± 1.1 71.5 71.4 71.4 42.8 94.1 77.7 71.0 97.8
olmOCR v0.1.75 (Anchored) [poznanski2025olmocr] 75.5 ± 1.0 74.9 71.2 71.0 42.2 94.5 78.3 73.3 98.3
MonkeyOCR-pro-3B [li2025monkeyocr] 75.8 ± 1.0 83.8 68.8 74.6 36.1 91.2 76.6 80.1 95.3
MinerU2.5 [niu2025mineru2] 77.5 ± 1.0 81.1 74.0 85.1 33.8 96.3 65.5 89.8 94.4
dots.ocr [dotsocr] 79.1 ± 1.0 82.1 64.2 88.3 40.9 94.1 82.4 81.2 99.5
@@
Table 4: Comprehensive evaluation of document parsing on olmOCR-Bench. Results are reported by olmOCR-Bench [poznanski2025olmocr] unless MinerU2.5 and Ours.
@@ 4.2.1 Text Recognition
Qwen2.5-VL-72B [bai2025qwen2] 0.054 0.023 0.061 0.084 0.195 0.032 0.056 0.118 0.040
MonkeyOCR-pro-3B [li2025monkeyocr] 0.058 0.021 0.064 0.096 0.116 0.023 0.058 0.124 0.052
MinerU2.5 [niu2025mineru2] 0.195 0.089 0.111 0.234 0.194 0.147 0.056 0.142 0.094
Dolphin [feng2025dolphin] 0.237 0.095 0.135 0.347 0.248 0.233 0.121 0.309 0.213
@@
Qwen2.5-VL-72B [bai2025qwen2] 0.405 0.056 0.389 0.165 0.194 0.758 0.164 0.220 0.021 0.181
Dolphin [feng2025dolphin] 0.682 0.699 0.912 0.691 0.709 0.832 0.818 0.549 0.037 0.309
MonkeyOCR-pro-3B [li2025monkeyocr] 0.601 0.182 0.921 0.449 0.876 0.909 0.896 0.387 0.036 0.262
MinerU2.5 [niu2025mineru2] 0.978 0.917 0.957 0.661 0.880 0.937 0.915 0.832 0.063 0.588
@@
Dolphin [feng2025dolphin] 0.236 0.145 0.074 0.025 0.095 0.218 0.113 0.183 0.092 0.190 0.202 0.225 0.230
MonkeyOCR-pro-3B [li2025monkeyocr] 0.253 0.071 0.048 0.023 0.295 0.529 0.144 0.165 0.063 0.086 0.110 0.184 0.263
Qwen2.5-VL-72B [bai2025qwen2] 0.188 0.047 0.037 0.018 0.100 0.387 0.122 0.186 0.034 0.090 0.041 0.134 0.220
MinerU2.5 [niu2025mineru2] 0.370 0.088 0.041 0.023 0.232 0.950 0.179 0.256 0.048 0.962 0.097 0.174 0.337
@@
This is a line and paragraph levels handwritten evaluation dataset designed for comprehensive handwriting recognition assessment. It contains 400 samples, evenly divided into four subsets of 100 images each. The dataset covers both real and synthetic handwriting in Chinese and English. Real samples are collected from established handwriting datasets such as CASIA-HWDB [liu2011casia], GNHK [lee2021gnhk], and BRUSH [kotani2020generating], while synthetic samples are generated to simulate diverse writing styles, character densities, and layouts. The benchmark aims to provide balanced and fine-grained evaluation for handwritten text recognition across different scripts and writing conditions.
@@
InternVL2.5-4B [InternVL] 0.197 0.240 0.661 0.741 0.674 0.754 0.655 0.734 0.406 0.473 0.652 0.687
MiniCPM-V2.6-8B [yao2024minicpm] 0.147 0.175 0.727 0.810 0.747 0.811 0.714 0.812 0.443 0.583 0.727 0.774
Qwen2-VL-7B [qwen2] 0.127 0.113 0.760 0.881 0.773 0.884 0.754 0.884 0.490 0.666 0.756 0.859
GOT [wei2024general] 0.616 0.402 0.283 0.568 0.309 0.618 0.273 0.544 0.151 0.295 0.255 0.492
PaddleOCR [cui2025paddleocr] 0.418 0.325 0.237 0.664 0.232 0.646 0.263 0.700 0.069 0.431 0.236 0.648
@@
Ocean-OCR [chen2025ocean] 0.145 0.106 0.774 0.885 0.780 0.912 0.782 0.862 0.532 0.736 0.772 0.885
MinerU2.5 [niu2025mineru2] 0.238 0.356 0.558 0.619 0.547 0.623 0.574 0.622 0.344 0.489 0.553 0.601
@@
Table 7: Comparison of performance on English(EN) and Chinese(ZH) OCR for handwritten recognition on Ocean-OCR-Bench. Results are reported by Ocean-OCR [chen2025ocean] unless MinerU2.5 and Ours.
@@
For table recognition, we utilize two benchmarks to validate the effectiveness of PaddleOCR-VL-0.9B based on TEDS [teds] and Edit Distance.
@@
dots.ocr [dotsocr] 0.8194 0.8442 0.1508
MinerU2.5 [niu2025mineru2] 0.9005 0.9539 0.0693
@@
MinerU2-VLM [MinerU2] 0.8286 0.8730 0.8757 0.9088
MonkeyOCR [li2025monkeyocr] 0.7396 0.7824 0.8174 0.8537
Nanonets-OCR-s [Nanonets-OCR-S] 0.7824 0.8190 0.8377 0.8692
OCRFlux-3B [OCRFlux2025] 0.7741 0.8071 0.8238 0.8617
Qwen2.5-VL-3B [bai2025qwen2] 0.7398 0.7765 0.8132 0.8701
Qwen2.5-VL-7B [bai2025qwen2] 0.7549 0.7926 0.8251 0.8819
Qwen2.5-VL-72B [bai2025qwen2] 0.7762 0.8361 0.843 0.8987
dots.ocr [dotsocr] 0.7547 0.7914 0.8047 0.8361
MinerU2.5 [niu2025mineru2] 0.8469 0.8955 0.8896 0.9239
@@
For formula recognition, we validate the effectiveness our model based on the Character Detection Matching (CDM) [cdm] metric on OmniDocBench-Formula-block and In-house-Formula datasets.
@@
dots.ocr [dotsocr] 0.4641 0.4868 0.4414
MinerU2-VLM [MinerU2] 0.8286 0.9616 0.6956
MonkeyOCR-pro-1.2B [li2025monkeyocr] 0.8531 0.9642 0.7419
MonkeyOCR-3B [li2025monkeyocr] 0.8621 0.9718 0.7524
Qwen2.5-VL-72B [bai2025qwen2] 0.8747 0.9574 0.7920
MinerU2.5 [niu2025mineru2] 0.9187 0.9751 0.8623
@@
Table 10: Comparison of OmniDocBench v1.5 Formula-block Performance. Due to dots.ocr [dotsocr] easily recognizing cropped formulas as images, the score is relatively low.
@@
dots.ocr [dotsocr] 0.6737 0.8066 0.5408
MinerU2-VLM [MinerU2] 0.9237 0.9764 0.8709
MonkeyOCR-pro-1.2B [li2025monkeyocr] 0.9537 0.9656 0.9417
MonkeyOCR-3B [li2025monkeyocr] 0.9566 0.9761 0.9371
Qwen2.5-VL-72B [bai2025qwen2] 0.9412 0.9519 0.9304
MinerU2.5 [niu2025mineru2] 0.9770 0.9832 0.9708
@@
Table 11: Comparison of In-house-Formula Performance. Due to dots.ocr [dotsocr] easily recognizing cropped formulas as images, the score is relatively low.
@@
For chart recognition, considering the limitations in dataset size, the imbalanced distribution of chart categories, and the poor annotation quality of publicly available test sets, we only utilize a in-house benchmark to validate the effectiveness of PaddleOCR-VL-0.9B based on the RMS-F1 [deplot] score metric. As shown in Table 12, the proposed PaddleOCR-VL not only outperforms expert OCR VLMs but also surpasses some 72B-level multimodal language models.
@@
TinyChart [zhang2024tinychart] 0.2159 0.4726 0.0876
GOT [wei2024general] 0.3160 0.1100 0.4190
OneChart [onechart] 0.3716 0.1384 0.4882
Qwen2.5-VL-3B [bai2025qwen2] 0.5942 0.5619 0.6103
Qwen2.5-VL-7B [bai2025qwen2] 0.6821 0.5876 0.7293
Qwen2.5-VL-72B [bai2025qwen2] 0.7300 0.6972 0.7464
PP-StructureV3 [cui2025paddleocr] 0.8060 0.7963 0.8109
@@ 4.3 Inference Performance
To improve the inference performance of PaddleOCR-VL, we introduce multi-threading asynchronous execution into the inference workflow. The process is divided into three main stages—data loading (e.g., rendering PDF pages as images), layout model processing, and VLM inference—each running in a separate thread. Data is transferred between adjacent stages via queues, enabling concurrent execution for higher efficiency. In particular, for VLM inference, batch processing is only triggered when either the number of items in the queue reaches a predefined threshold or the waiting time for queued data exceeds a specified limit. This design allows blocks across different pages to be aggregated and processed together, thereby maximizing parallelism, especially when handling large volumes of files. We further deploy PaddleOCR-VL-0.9B on high-throughput inference and serving engines [kwon2023efficient, zheng2024sglang, PaddlePaddle_FastDeploy], tuning parameters like max-num-batched-tokens and gpu-memory-utilization to balance inference throughput with GPU memory consumption.
@@
MonkeyOCR-pro-1.2B [li2025monkeyocr] vLLM (v0.10.2) 1456.4 0.6730 1120.3
dots.ocr [dotsocr] vLLM (v0.10.2) 2784.6 0.3522 532.9
MinerU2.5 [niu2025mineru2] vLLM (v0.10.2) 927.3 1.0574 1647.9
@@ A.2 Table
As shown in Figure A2, we constructed a large-scale dataset of over $5$ million high-quality image-table pairs. Our dataset construction employs three key strategies: automatic data annotation, potential annotation mining, and high-quality data synthesis. For coding efficiency, we adopt OTSL [lysak2023optimized] as the model’s target format instead of conventional HTML. The main dataset construction process is as follows:
@@
    Automatic Data Annotation: To enhance the performance of PaddleOCR-VL in table recognition, we built a large-scale, diverse dataset covering various languages, border styles, and table types. Tables are first located using PP-StructureV3 [cui2025paddleocr]. For unlabeled images, we employed a multi-stage annotation pipeline: ERNIE-4.5-VL [ernie2025technicalreport] first generates pseudo-labels, which are then validated by a ERNIE-4.5-VL-28B-A3B [ernie2025technicalreport] as discriminative model. Rejected annotations are refined using DianJin-OCR-R1 [chen2025dianjin] (for tools, we use ERNIE-4.5-VL and PP-StructureV3 [cui2025paddleocr]). Finally, all annotations undergo rigorous rule-based verification, including n-gram analysis and HTML validation, to ensure only high-quality samples are used for training.